In a high school of 800 students, 45% of them are boys. If 66 2/3% of them smoke, how many boys in the school smoke?
( I really need help.. especially with fractions, decimals and percent)

Answer by stanbon(75887)

In a high school of 800 students, 45% of them are boys. If 66 2/3% of them smoke, how many boys in the school smoke?
-------
45% = 45/100 = 0.45
----
66 2/3% = (66 2/3)/100 = 0.66666....
==============
# of boys in the school is 0.45*800 = 360
----
Cheers,
Stan H.
===========